The perimeter of a rectangle is 196 inches. The length exceeds the width by 44 inches. Find the length and width???
Answer by rapaljer(4671)

Let x = width
x+44 = length

2W + 2 L = P
2(x)+2(x+44)=196
2x+2x +88=196
4x=196-88
4x=108
x=27 width
x+44=27+44=71 length

Check:
2*27+2*71
54+142
196
It checks!!

For additional help on WORD PROBLEMS, please see my own website for a LOT of FREE resources. Click on my tutorname "rapaljer" anywhere in algebra.com. Once you find my Homepage, look for "Basic, Intermediate, and College Algebra: One Step at a Time". There you will find my own user-friendly explanation for just about all the topics in algebra. Choose "Basic Algebra", then look in "Chapter 1" for "Sections 1.09 and 1.10".

Many of my exercises in this section are solved in the "MATH IN LIVING COLOR" sections that go with this. There you will find an explanation that will probably be a lot easier to understand than your own expensive textbooks!! My materials are all free.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus